OpenStudy (anonymous):

sinx*tanx=1-cos^2x/cosx

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf sin~x~tan~x~=~\frac{1-cos^{2}x}{cos~x} }\) like this ?

OpenStudy (anonymous):

\[\sin(x)\tan(x)=\frac{ \sin(x) }{ \cos(x) }\] \[\[\sin(x)\tan(x)=\tan(x)\] \[\sin(x)=1\] \[x=\sin ^{-1}(1)=90\]

OpenStudy (solomonzelman):

No no no !!!! \(\LARGE\color{blue}{ \bf sin~x~tan~x~=~\frac{1-cos^{2}x}{cos~x} }\) \(\LARGE\color{blue}{ \bf sin~x~tan~x~=~\frac{sin^{2}x}{cos~x} }\) \(\LARGE\color{blue}{ \bf sin~x~tan~x~=sin~x~~\frac{sinx}{cos~x} }\) \(\LARGE\color{blue}{ \bf sin~x~tan~x~=sin~x~tan~x~ }\)

OpenStudy (anonymous):

Ohh, that's true I forgot it is equal to sin^2(x) xD

OpenStudy (solomonzelman):

Yeah :)